# CaRBM: A Fixed-Depth Quantum Algorithm with Partial Correction for Thermal State Preparation

### 18 Mar 2026 v1

@@ I Introduction
Alternative approaches that avoid these challenges are based on imaginary time evolution (ITE). These methods — also known as purification — are frequently employed in classical computing methods, like tensor networks, as well as in quantum algorithms [12, 13, 14]. The main drawback of such quantum algorithms is that the operator needed for ITE is not unitary, which makes it unimplementable on a quantum device without further treatment like block encoding or approximating the effects of the non-unitary operations by unitaries [15, 16, 17, 18, 19, 20].
@@
The second approach is to embed the non-unitary operator inside a larger unitary matrix, then recover the effects of the non-unitary operation by post-selection. This is known as block-encoding [15, 16, 17, 18], and it is our method of choice in this paper. We employ a block-encoding scheme — based on a neural network architecture known as restricted Boltzmann machine (RBM) — that implements the non-unitary action on the physical (visible) qubits as part of a larger unitary operator by expanding the system with additional ancillary (hidden) qubits [18] (see Fig. 1). In the original algorithm, operators of the form $e^{-\kappa\sigma}$, where $\sigma$ is a Pauli string, are expressed as a sum of unitary operators, which can be realized as a quantum circuit with an ancillary qubit. The operation is then accepted with some acceptance probability depending on the outcome of the measurement on the ancilla.
@@ III.2 Thermal phase diagram of relativistic fermions in $1+1d$
In our simulations, we set $N=2$ and $L=4$
and calculate the two-point correlation function $\sum_{i}\langle\bar{\psi}_{i}\psi_{0}\rangle\sim\sum_{i}\langle Z_{i}Z_{0}\rangle$. Again, we start with an ${\rm TFD}(0)$ and obtain the desired ${\rm TFD}(\beta)$ via imaginary time evolution using $H_{GN}$.
We find that the phase diagram obtained with our method which can be seen in Fig. 5 is very similar to the known results, even for small lattice sizes and number of flavors [45, 46, 47]. As expected, we see a symmetry-broken phase with a finite condensate for large $\beta$ and small $\mu$ indicated as the bright region in the graph, and a symmetric phase for large $\mu$ indicated as the black region. However, we don’t see a sign of an inhomogeneous phase due to the small lattice volumes and number of flavors. The effect of the correction scheme is also apparent in these phase diagram plots. In the top left plot, we clearly see an unphysical phase on the top right. After implementing the correction scheme, we see that this region disappears, and we obtain the expected phase diagram.
@@ Acknowledgments
AFK acknowledges financial support from the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research under DE-SC0025623 and GCT and OA from the National Science Foundation under award No. PHY-2325080: PIF: Software-Tailored Architecture for Quantum Co-Design. ER is supported by the U.S. Department of Energy (DOE) under Contract No. DE-AC02-05CH11231, through the National Energy Research Scientific Computing Center (NERSC), an Office of Science User Facility located at Lawrence Berkeley National Laboratory.

### 23 Mar 2026 v2

@@ I Introduction
Alternative approaches that avoid these challenges are based on imaginary time evolution (ITE). These methods — also known as purification — are frequently employed in classical computing methods, like tensor networks, as well as in quantum algorithms [12, 13, 14]. The main drawback of such quantum algorithms is that the operator needed for ITE is not unitary, which makes it unimplementable on a quantum device without further treatment like block encoding or approximating the effects of the non-unitary operations by unitaries [15, 16, 17, 18, 19, 20, 21].
@@
The second approach is to embed the non-unitary operator inside a larger unitary matrix, then recover the effects of the non-unitary operation by post-selection. This is known as block-encoding [16, 17, 24, 18, 19], and it is our method of choice in this paper. We employ a block-encoding scheme — based on a neural network architecture known as restricted Boltzmann machine (RBM) — that implements the non-unitary action on the physical (visible) qubits as part of a larger unitary operator by expanding the system with additional ancillary (hidden) qubits [18] (see Fig. 1). In the original algorithm, operators of the form $e^{-\kappa\sigma}$, where $\sigma$ is a Pauli string, are expressed as a sum of unitary operators, which can be realized as a quantum circuit with an ancillary qubit. The operation is then accepted with some acceptance probability depending on the outcome of the measurement on the ancilla.
@@ III.2 Thermal phase diagram of relativistic fermions in $1+1d$
In our simulations, we set $N=2$ and $L=4$ and calculate the two-point correlation function $\sum_{i}\langle\bar{\psi}_{i}\psi_{0}\rangle\sim\sum_{i}\langle Z_{i}Z_{0}\rangle$. Again, we start with the infinite temperature state and obtain the desired state via imaginary time evolution using $e^{-\beta H_{GN}}$. We find that the phase diagram obtained with our method which can be seen in Fig. 5 is very similar to the known results, even for small lattice sizes and number of flavors [46, 47, 48]. As expected, we see a symmetry-broken phase with a finite condensate for large $\beta$ and small $\mu$ indicated as the bright region in the graph, and a symmetric phase for large $\mu$ indicated as the black region. However, we don’t see a sign of an inhomogeneous phase due to the small lattice volumes and number of flavors. The effect of the correction scheme is also apparent in these phase diagram plots. In the top left plot, we clearly see an unphysical phase on the top right. After implementing the correction scheme, we see that this region disappears, and we obtain the expected phase diagram.
@@ Acknowledgments
AFK acknowledges financial support from the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research under DE-SC0025623 and GCT and OA from the National Science Foundation under award No. PHY-2325080: PIF: Software-Tailored Architecture for Quantum Co-Design. ER is supported by the U.S. Department of Energy (DOE) under Contract No. DE-AC02-05CH11231, through the National Energy Research Scientific Computing Center (NERSC), an Office of Science User Facility located at Lawrence Berkeley National Laboratory. EJR was supported by the National Science Foundation under cooperative agreement 2020275 during partial completion of this work. EJR is supported by the U.S. Department of Energy through the Los Alamos National Laboratory. The research presented in this article was partially supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project numbers 20251163PRD3 and 20260043DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).